(1) Convert the decimal numbers +61 and +27 to their 8-bit 2s complement representations. Then perform the binary equivalent of (b) (27)+ (+61); and Convert the answers back to decimal and verify that they are correct.
Answer #1

If the decimal number is positive, the two’s-complement number is the binary equivalent of the decimal number. If the decimal

Determine the 2s complement of (270). -2710 0111001002 + 1s complement number + 1 Add 1 to the ls complement number -2710

Determine the 2s complement of (610). -6110 0110000102 → ls complement number + 1 Add 1 to the ls complement number -61,0

Therefore, [(ox 2?)+(0x2°)+(1x 2°)+(0x24) ) 001000102 = | +(0x2°)+(0x22)+(1x2-)+0x2°) = 0+0+32 +0+0+0+2+0 = 34 The equivalent

001000 Verify the above value is correct or not. on f(0x2?)+(0x 20)+(1x 2°) +(0x 2^) | +(0x2?)+(0x22)+(1x2-)+(0x 2°) = 0+0+32

Verify the above value is correct or not. The 2’s complement of (10101000) is, 01010111, → ls complement number + 1 Add 1 to

> This was really really helpful. Thank you

Kemo Robson Sun, Dec 5, 2021 9:20 PM